how to solve 5/12 of 60 1/8 of 2 3/4 of 80 4/7 of 77 1/2 of 1/8
Answer by checkley77(12844)

5/12 of 60
5/12*60=300/12=25 ans.
1/8 of 2
2/8=1/4 ans.
3/4 of 80
3/4*80=240/4=60 ans.
4/7 of 77
4*77/7=4*11=44 ans.
1/2 of 1/8
1/2*1/8=1/16 ans.